# Improving Algorithmic Efficiency using Cryptography

### v1

@@ 6 Summary and Discussion
The main contribution of this work is conceptual: we propose speeding up classical algorithms by replacing a truly random object with a pseudorandom one that contains a trapdoor — of a form that allows an equivalent yet more efficient computation. We raise the question of whether this idea can be extended beyond random matrices to other objects.
Another important direction is to further support or refute Conjecture 2.3, for instance, by showing it is implied by more standard cryptographic assumptions, or by showing that natural attack strategies are bound to fail.
Acknowledgments:
OZ’s research is supported in part by the Israel Science Foundation, Grant No. 1593/24, and in part by the Blavatnik Family foundation. VV’s research is supported in part by DARPA under Agreement Number HR00112020023, NSF CNS-2154149, a Simons Investigator award and a Thornton Family Faculty Research Innovation Fellowship from MIT.

### v2

@@ 7 Summary, Discussion and Open Problems
The main contribution of this work is conceptual: we propose accelerating classical algorithms by replacing a truly random object with a carefully structured pseudorandom object that comes with a (hidden) trapdoor that enables an equivalent but more efficient computation. We believe this idea opens a much broader direction of inquiry, and to that end, we raise the question of whether similar constructions can be applied beyond random matrices to other classes of random objects.
This suggests a general framework: we are interested in meaningful sets of actions $\mathcal{G}$ that act on a set of objects $\mathcal{X}$. In the exemplary scenarios explored in this paper, $\mathcal{G}$ is the set of all linear transformations (represented as matrices) and $\mathcal{X}$ is either the set of all vectors, or the set of all matrices themselves. We ask: are there other interesting distributions over actions $\mathcal{D}(\mathcal{G})$ for which the following holds?
  • Applying a random action $g\sim\mathcal{D}(\mathcal{G})$ (chosen from the prescribed distribution) to an object $x\in\mathcal{X}$ is computationally expensive.
  • There exists a pseudorandom distribution $\tilde{\mathcal{D}}$ of actions, indistinguishable from the original, for which performing the same action is significantly faster, given a trapdoor $t_{g}$ for $g\sim\mathcal{D}(\mathcal{G})$.
Beyond matrix-vector and matrix-matrix multiplication, other potential examples in the matrix setting include operations such as repeated exponentiation, which can also be viewed as taking multiple steps in a random walk. More broadly, other random structures that may admit similar trapdoor constructions include random graphs, random sets, and random polynomials. We leave the exploration of these objects and their potential applications to future work.
We also leave open the problem of further supporting or refuting Conjecture 5.1. This could involve proving that the conjecture follows from standard cryptographic assumptions, or alternatively, demonstrating that plausible classes of attack strategies, e.g. attacks implementable via low-degree polynomials, are provably ineffective.
Finally, several problems remain open in connection to the (conditional) worst-case to average-case reductions for linear algebraic operations. In particular, can errors be corrected for the problems in addition to matrix multiplication? What is the lowest success probability from which an average-case algorithm for the determinant can be used for worst-case inputs? And for which other operations such reductions can be obtained?
Acknowledgements.
The authors would like to thank Mark Braverman and Stephen Newman for sending us a draft of their work [BN25].